Solve by the elimination method. 2x - 4y = 3
2x - 4y = 4
What is the solution of the system?

Answer by checkley77(12844)

2x - 4y = 3
2x - 4y = 4 subtract.
------------------------
0x+0y=-1 no unique solution for x or y.